number of four digit numbers in which at least one digit occurs more than once is
 
 

Grade:11

1 Answers

Deepak Kumar Shringi
5 years ago
total four digit numbers=9*10*10*10=9000
in which non repetitng =9*9*8*7=4536
remaining = 9000-4536=3464